find all zeros of the polynomial function. g(x)=x^3-2x^2-x+2
Found 2 solutions by jim_thompson5910, Alan3354:
Answer by jim_thompson5910(35256)

Hint:

g(x)=x^3-2x^2-x+2

0=x^3-2x^2-x+2

x^3-2x^2-x+2 = 0

(x^3-2x^2)+(-x+2) = 0

x^2(x-2)-1(x-2) = 0

(x^2 - 1)(x-2) = 0

(x-1)(x+1)(x-2) = 0

Answer by Alan3354(69443)

find all zeros of the polynomial function. g(x)=x^3-2x^2-x+2
--------------
Try x = 1 & x = 2, plus and minus first.